OpenStudy (anonymous):

4 log x-6 log (x+2) write the expression as a single logarithm. a.24 log x/x+2 b.log x^4 (x+2)^6 c.logx(x+2)^24 d.none

OpenStudy (inkyvoyd):

Do you remember your logarithm laws?

OpenStudy (anonymous):

4logx = log(x^4) -6log(x+2) = log((x+2)^-6) so 4logx -6log(x+2) = log(x^4)+log((x+2)^-6) = log(x^4*(x+2)^-6) = log (x^4/(x+2)^6)

OpenStudy (inkyvoyd):

that works too ._.